OpenStudy (anonymous):

Can someone pls help explain how they came up with this equation: 212F=__C C=5/9 C-32 C=5/9(212)-32 C=5/9(180) C=100 So 212F = 100C any help thnks.

OpenStudy (campbell_st):

I think it works this way \[C = \frac{5}{9} \times (212 - 32)\] order of operation says brackets 1st \[C = \frac{5}{9} \times 180\] \[C = \frac{900}{9} \] C = 100

OpenStudy (anonymous):

ok but how did you get 900

OpenStudy (anonymous):

what made you go from 180 to 900

OpenStudy (campbell_st):

180 x 5 = 900

OpenStudy (anonymous):

k

OpenStudy (campbell_st):

5/9 x 180 = (5 x 180)/9

OpenStudy (anonymous):

did you set 5/9 x 180 up as a fraction that needed to be reduced

OpenStudy (anonymous):

5/9 x180/1

OpenStudy (campbell_st):

I did so you could see the steps... normally I'd type it into a calculator as I've written it at the start 5/9 x(212 - 180)

OpenStudy (campbell_st):

no you don't need both as fractions

OpenStudy (anonymous):

im a lil confused im studying some math problems for my nxt school yr im an 8th grader so dont be mad if i keep asking questions...lol thnks for the help

OpenStudy (campbell_st):

thats ok...

OpenStudy (anonymous):

im not good at this so thats y im studying these type of math quest.